OpenStudy (anonymous):

find f. f'(x)=6 cos t + sec^2t, −π/2 < t < π/2, f(π/3) = 1

OpenStudy (anonymous):

i got 6 sin(t)+tan(t)+1-2sqrt(3)

OpenStudy (anonymous):

any ideas

OpenStudy (dumbcow):

to find f, you integrate f' looks good, i think the constant is a little off though sin(pi/3) = sqrt3/2 tan(pi/3) = sqrt3 --> c = 1-4sqrt3

OpenStudy (anonymous):

omg i missed that

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

no prob :)